OpenStudy (anonymous):

Bradley invested an average of $550 per month since age 49 in various securities for his retirement savings. His investments averaged a 7% annual rate of return until he retired at age 73. Given the same monthly investment and rate of return, how much more would Bradley have in his retirement savings had he started investing at age 40? Assume monthly compounding.

OpenStudy (anonymous):

$440,088.67 $63,558.00 $1,344,425.69 $282,480.00

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

you will use the FVOA formula (FVOA = future value of annuity)

jimthompson5910 (jim_thompson5910):

find the FVOA for when he starts at age 49 and goes til age 73 then find the FVOA for when he starts at age 40 and goes til age 73 and take the difference of the two

jimthompson5910 (jim_thompson5910):

oh sry, FVOA = future value of ordinary annuity but same steps will be taken to get the answer

OpenStudy (anonymous):

that formula is the FVOA= C((1+i)^nt) / i

jimthompson5910 (jim_thompson5910):

yes, C is the monthly contribution so C = 550 i = r/n = 0.07/12 = 0.00583 (compounding monthly) n = 12 (compounding monthly) t = number of years

OpenStudy (anonymous):

t = 33?

OpenStudy (anonymous):

or does t =9?

OpenStudy (anonymous):

neither of those are right :(

jimthompson5910 (jim_thompson5910):

when you go from age 40 to age 73, the value of t is 33 (73 - 40 = 33) when you go from age 49 to age 73, the value of t is 24 (73 - 49 = 24)

jimthompson5910 (jim_thompson5910):

so you need to find FVOA when t = 33 and when t = 24 then subtract the FVOA values

OpenStudy (anonymous):

ohh! okay

OpenStudy (anonymous):

i got 439583. but it's really close to one of my answers

jimthompson5910 (jim_thompson5910):

Try it again. The answer I get is one of the answer choices.

jimthompson5910 (jim_thompson5910):

btw, the formula you are working with is \[\Large FVOA = C*\frac{(1+i)^{nt}-1}{i}\]

jimthompson5910 (jim_thompson5910):

You said before that it was " FVOA= C((1+i)^nt) / i" but there should be a -1 up top and I missed that as well. But it's fixed now.

OpenStudy (anonymous):

thats what it was haha i forgot the -1 oops

OpenStudy (anonymous):

i got 439883?

jimthompson5910 (jim_thompson5910):

when t = 33, you should get FVOA = 849234.96

jimthompson5910 (jim_thompson5910):

check your work again

OpenStudy (anonymous):

for that i got 848482. i probably put in a wrong number or something.

OpenStudy (anonymous):

and for the t=24 i got 408,899

jimthompson5910 (jim_thompson5910):

what are you typing in for t = 33?

OpenStudy (anonymous):

550((1+.00583)^396 -1) / .00583

jimthompson5910 (jim_thompson5910):

maybe you need to use more decimal digits? If you do, 550((1+.005833333)^396 -1) / .005833333 you should get 849234.88

jimthompson5910 (jim_thompson5910):

it's probably better to replace .005833333 with 0.07/12

jimthompson5910 (jim_thompson5910):

doing that gives 550((1+0.07/12)^396 -1) / (0.07/12) = 849,234.96

OpenStudy (anonymous):

yeah that worked

jimthompson5910 (jim_thompson5910):

do the same for t = 24, what do you get?

OpenStudy (anonymous):

408490

jimthompson5910 (jim_thompson5910):

I'm getting 409146.29 for t = 24

jimthompson5910 (jim_thompson5910):

hmm not sure how you got that, did you do 550((1+0.07/12)^288-1) / (0.07/12) or no?

jimthompson5910 (jim_thompson5910):

what do you get when you compute 550((1+0.07/12)^288-1) / (0.07/12)